OpenStudy (anonymous):

MEDAL AND FAN! Second-degree polynomial is given by f : f(x)=0.5x^2-4x+6 Determine the peak for the corresponding parabola and any intersections with the x-axis. Determine then value the amount of f (Remember to explain how to get to the answer).

OpenStudy (anonymous):

Peak as in highest point? or lowest point?

OpenStudy (whpalmer4):

Find the vertex of the parabola. Do you know how to do that?

OpenStudy (anonymous):

@OrthodoxMan peak meaning top point, which should be the highest point ...

OpenStudy (anonymous):

@whpalmer4 no, can you help?

OpenStudy (whpalmer4):

The formula for this parabola is \[y = f(x) = 0.5x^2-4x+6\]which corresponds to the form\[y = ax^2+bx+c\]If \(a>0\) the parabola opens upward, and the vertex is at the bottom. Are you sure that you have the formula correct? Your question somewhat suggests that the vertex is at the highest point, but with this equation, it will be at the lowest point.

OpenStudy (anonymous):

to find the vertex, use the following steps 1. Group x terms in a bracket 2. Factor out coefficient of x^2 3. Divide the other x term by 2 then square it. Take that number, add it and subtract it in the bracket. 4. The subtracted number is to be multiplied by 0.5 and subtracted out of the bracket with the number outside. 5. Factor whatever you have left. Ex. 0.5(x+____)^2 +/- ____

OpenStudy (whpalmer4):

In any case, if the equation is in the form \[y = ax^2+bx+c\] the x-coordinate of the vertex will be at \[x=-\frac{b}{2a}\]and the y coordinate can be found by plugging that value of \(x\) into \(f(x)\)

OpenStudy (whpalmer4):

@orthodoxman's approach works as well, but is considerably more work :-)

OpenStudy (anonymous):

True true :P

OpenStudy (whpalmer4):

Now, if you happen to already have the roots (because it is written in factored form, for example), then it makes getting the vertex quite easy, as the vertex is exactly halfway between the roots.

OpenStudy (anonymous):

@whpalmer4, well instead of 0.5 it's says 1/2. But that's about it... when I use the x=-b/2a, I get 4... What do I do now?

OpenStudy (anonymous):

@whpalmer4 Do i do f(4)=0.5(4)^2-4(4)+6?

OpenStudy (whpalmer4):

1/2 = 0.5, of course

OpenStudy (whpalmer4):

Yes, that's the value of y at the vertex

OpenStudy (anonymous):

So f(4) = -2 ? @whpalmer4

OpenStudy (whpalmer4):

It does when I compute it...

OpenStudy (anonymous):

@whpalmer4, what do i do now? :P

OpenStudy (whpalmer4):

Now you need to find the intersections with the x-axis. Those are the values of x such that f(x) = 0.

OpenStudy (anonymous):

@whpalmer4, is the "top point" then (4,-2) ?

OpenStudy (whpalmer4):

well, the vertex is at (4,-2), but it isn't the top point, it's the bottom point. That's why I asked about whether your equation was possibly incorrect. With a negative sign in front of the leading term, we would have a parabola that opens downward, and that would have something that could reasonably be described as a "top point"

OpenStudy (anonymous):

But that's all it says. I'll just write that and then see if i get it right. But that would be the point, right? @whpalmer4

OpenStudy (whpalmer4):

You do still have to find the points where the parabola crosses the x-axis...

OpenStudy (anonymous):

@whpalmer4 yeah, the intersections. How do I do that? Can you help?

OpenStudy (whpalmer4):

@orthodoxman actually showed you a procedure for doing so. You can also use the formula for the solutions of a quadratic: if \[y =ax^2+bx+c,\, a\ne 0\]then the solutions are \[x = \frac{-b\pm\sqrt{b^2-4ac}}{2a}\]

OpenStudy (anonymous):

@whpalmer4 , I know y should 0. Because it's the intersections that's on the x-axis we need to find... right?

OpenStudy (whpalmer4):

yes, y =0 at those points.

OpenStudy (anonymous):

@whpalmer4 Thanks! I'll do that first!

OpenStudy (anonymous):

@whpalmer4 I got x=1 or x=1/3 .... do you know if that's right?

OpenStudy (whpalmer4):

plug in the values in the original equation. do you get y = 0 if so? \[y = \frac{1}{2}x^2-4x+6\] \[y = \frac{1}{2}(1)^2-4(1)+6\]\[y=\frac{1}{2}+2\]Nope, not correct.

OpenStudy (whpalmer4):

Also, the two solutions are symmetrical about the vertex. If the vertex is at x = 4, one solution is at x-a, the other at x + a, for some value of a.

OpenStudy (anonymous):

Find 2 x-intercepts: y = x^2/2 - 4x + 6 = 0 = x^2 - 8x + 12 = 0. Both roots are positive. Compose factor pairs of c = 12. Proceeding: (1, 12)(2, 6). This last sum is 2 + 6 = 8 = -b. Then, the 2 real root are 2 and 6. Min at x = -b/2a = 4/1 = 4 Ymin. = 8 - 16 + 6 = -2.

OpenStudy (anonymous):

@whpalmer4 i have no idea what i'm doing now

OpenStudy (anonymous):

@thu1935 yeah! I've done that one :-)

OpenStudy (whpalmer4):

Why don't you show me the work you did with the quadratic formula that gave you the answers of 1 or 1/3? Let's find out where you made the mistake

OpenStudy (whpalmer4):

First, what are you using for a, b, c?

OpenStudy (anonymous):

I'm using a= 1/2, b=-4 and c=6

OpenStudy (anonymous):

@whpalmer4

OpenStudy (anonymous):

@whpalmer4, you don't know what I did wrong ?

OpenStudy (whpalmer4):

the denominator is \(2a\) not \(2ac\)

OpenStudy (anonymous):

@whpalmer4 oh, yeah. Crap.

OpenStudy (anonymous):

So it's 6 and 2?

OpenStudy (anonymous):

Yup! it's right! @whpalmer4

OpenStudy (whpalmer4):

And 6 and 2 are symmetrically located about 4. Graph looks like this:

OpenStudy (anonymous):

Anyway. Thanks a bunch, @whpalmer4 and @OrthodoxMan!

OpenStudy (whpalmer4):

You're welcome!